A farmer has 200 feet of fencing. Using a barn as one side, the farmer wants to create a rectangular pen with three areas using the fencing. What are the dimensions of the pen that give the farmer the greatest enclosed area?
Answer by richwmiller(17219)

200+b is the perimeter.
100+(100+50+50)=
100*50=5000
100*50 is the pen
using 100 ft of barn
x+.5x+.5x=200
2x=200
x=100
.5x=50